### v2

# LSAST: Enhancing Cybersecurity through LLM-supported Static Application Security Testing

Abstract
The current cybersecurity landscape is increasingly complex, with traditional Static Application Security Testing (SAST) tools struggling to capture complex and emerging vulnerabilities due to their reliance on rule-based matching. Meanwhile, Large Language Models (LLMs) have demonstrated powerful code analysis capabilities, but their static training data and privacy risks limit their effectiveness. To overcome the limitations of both approaches, we propose LSAST, a novel approach that integrates LLMs with SAST scanners to enhance vulnerability detection. LSAST leverages a locally hostable LLM, combined with a state-of-the-art knowledge retrieval system, to provide up-to-date vulnerability insights without compromising data privacy. We set a new benchmark for static vulnerability analysis, offering a robust, privacy-conscious solution that bridges the gap between traditional scanners and advanced AI-driven analysis. Our evaluation demonstrates that incorporating SAST results into LLM analysis significantly improves detection accuracy, identifying vulnerabilities missed by conventional methods.
Index Terms:
Vulnerability Scanning; Large Language Model
I Introduction
With the global loss of cybercrime projected to increase from \$8.15 trillion in 2023 to \$13.82 trillion in 2028 [1] and hackers launching an average of 26,000 attacks daily—one every three seconds [2]—cybersecurity is more critical than ever in the digital age. As the IT landscape grows, cyber threats have evolved from simple attacks to sophisticated and persistent threats, necessitating the implementation of comprehensive cybersecurity measures.
One widely adopted measure is security scanning[3, 4] to find out the potential vulnerabilities [5, 6], which plays a crucial role in identifying potential risks within applications and networks. However, cybersecurity threats are continuously evolving, and new vulnerabilities emerge rapidly, making it increasingly difficult for traditional security detection tools to keep pace. Static Application Security Testing (SAST)[7], a conventional static analysis tool, is commonly used for code security testing, but its rule-based pattern matching approach limits its effectiveness in detecting complex vulnerabilities. Large Language Models (LLMs) [8, 9] have demonstrated significant potential in code analysis, particularly with their abilities to understand complex contexts and detect diverse vulnerabilities. By retrieving the latest vulnerability reports, LLMs can learn new types of vulnerabilities in real-time without being constrained by static training data[10, 11, 12]. Unlike traditional pattern matching and fuzzing techniques, LLMs can identify vulnerabilities that may be overlooked by conventional scanners[13, 14, 15].
However, since LLMs’ training data is not updated after the initial training, they struggle to promptly recognize newly emerging security threats. This limitation poses substantial security risks when using LLMs for vulnerability scanning, as effective scanning depends on accessing to up-to-date information on the latest vulnerabilities to adequately address emerging threats.
Additionally, using LLMs for vulnerability scanning introduces further challenges, particularly when code is transmitted to third-party LLM providers via API endpoints. This practice can expose sensitive code, leading to potential privacy and security risks, including unauthorized access by malicious actors. Moreover, there is a risk of the code being repurposed for future model training, making it vulnerable to training data extraction attacks. Addressing these concerns is essential to prevent the exploitation of potential vulnerabilities in the code.
In this paper, we aim to tackle these challenges head-on. We
  • enhance LLMs’ capabilities in vulnerability scanning by integrating static code scanning results with LLMs, allowing them to detect vulnerabilities that would be missed by traditional methods alone. This demonstrates the powerful synergy achieved by combining these technologies.
  • address the limitation of LLMs in accessing up-to-date vulnerability information by introducing novel approaches for knowledge retrieval systems. Our combined and abstraction-based approaches outperform existing methods and set new standards for building and enhancing knowledge retrieval systems for static vulnerability scanning.
  • demonstrate that knowledge retrieval systems can be counterproductive if not carefully designed, as irrelevant reports can distract the LLM’s focus.
  • mitigate privacy and security risks by utilizing open-source LLMs, showing that LLMs can be leveraged for vulnerability scanning without compromising these critical areas.
  • validate in real-world scenarios that LSAST scanners can identify vulnerabilities overlooked by traditional scanners, establishing their importance in comprehensive security scanning processes.
The rest of the paper is organized as follows. Section II presents the background of security scanners and vulnerability information. Section III reviews the related work on security testing tools, LLMs in vulnerability detection and Knowledge retrieval. Section IV gives the LSATA and details the optimization based on SAST. Section V presents our experimental results and findings. Finally, Section VI concludes the paper.
II Background
In this Section, we will first introduce different kinds of existing security scanners, followed by Retrieval-Augmented Generation (RAG).
II-A Different kinds of security scanners
Predominantly, security scanning can be classified into two categories: DAST (Dynamic Application Security Testing) and SAST (Static Application Security Testing)[16]. DAST scanners treat the system or application as a black-box, meaning they operate without knowledge of the internal workings. They identify weaknesses by interacting with the system or application during runtime. By sending specially crafted requests and analyzing the responses, DAST scanners mimic threat actors attempting various types of attacks, such as SQL Injection, Cross-Site Scripting, and others. SAST scanners, also known as Source Code Analysis Tools, examine the source code without executing the application. They search for specific patterns in the code that are known to be vulnerable. Both types of scanners complement each other, as they can detect vulnerabilities that the other type might miss. Thereby it is crucial for stakeholders to incorporate both types of scanners into their infrastructure.
II-B Retrieval-Augmented Generation
As the cybersecurity landscape evolves rapidly, staying updated with the latest vulnerabilities and attack vectors is crucial for effective defense. New security threats, vulnerabilities, and coding errors are discovered frequently, and traditional approaches to static analysis struggle to keep pace with these emerging risks. LLMs, while powerful in detecting logic flaws and complex vulnerabilities, are limited by their static training data, which often does not include the most recent security developments.
To address this limitation, we leverage Retrieval-Augmented Generation (RAG), a technique that dynamically enhances LLM capabilities by integrating relevant external knowledge during the response generation process. RAG bridges the gap between static training data and the need for real-time awareness of the latest vulnerabilities. In RAG, external knowledge is stored as vector representations within specialized databases. During the LLM’s processing, similarity searches are conducted within these databases to retrieve the most relevant information. This approach allows the LLM to draw upon a continuously updated knowledge base, incorporating fresh insights and vulnerability reports that are not part of the original training data.
III Related Work
This section reviews existing research on vulnerability detection methods, highlighting the limitations of traditional approaches, the emerging role of LLMs, and the integration of knowledge retrieval systems.
III-A Limitations of Traditional Methods
Traditional static and dynamic application security testing tools, such as SAST and DAST, have long been the cornerstone of vulnerability detection. However, several studies have identified significant limitations in their effectiveness, particularly when dealing with complex and context-specific vulnerabilities. For example, SAST tools often rely on predefined patterns and rules, which can result in high false-positive rates and an inability to detect novel or context-dependent vulnerabilities. This issue is underscored in [17] and [18]. Attempts to integrate machine learning or heuristic-based approaches into SAST and DAST have been explored, yet they often lack the depth of understanding needed for sophisticated vulnerability detection. Li et al. [19] recognize the potential of combining LLMs with traditional static analysis but focus narrowly on specific bugs, such as Use-Before-Initialization (UBI) in Linux kernels, and do not fully leverage specific static analysis tools or support a wide range of programming languages. These gaps have led researchers to explore the use of LLMs as a way to enhance the effectiveness of traditional static analysis tools.
III-B Other Aspect of Vulnerability
Prior work has explored various aspects of security and privacy vulnerabilities in machine learning models. The study in [6] addresses model inversion attacks (MIAI), proposing a novel black-box approach to infer sensitive information about the training data while highlighting the disparate vulnerability impact on specific demographic groups. However, it focuses solely on privacy risks in machine learning models and does not address static code analysis or vulnerability detection in software applications. Meanwhile, [5] investigates the security implications of deploying efficient, on-device deep learning models, showing that these models can increase adversarial vulnerabilities in their full-scale counterparts. Although it introduces a defense mechanism to mitigate these risks, the work is limited to adversarial attacks and does not extend to static vulnerability detection. In contrast, our proposed LSAST framework addresses these limitations by integrating Large Language Models (LLMs) with traditional Static Application Security Testing (SAST) tools, enhancing vulnerability detection in software applications while maintaining data privacy and providing up-to-date insights, which are not covered in these previous works.
III-C LLMs in Vulnerability Detection
LLMs have emerged as a powerful tool in the field of vulnerability detection, offering a significant upgrade over traditional scanners due to their ability to understand code in a more context-aware manner. Studies such as [20] and [21] have demonstrated that LLMs can effectively identify vulnerabilities that static tools often miss, thanks to their deep learning-based understanding of code structure and semantics. Despite these advantages, LLMs face their own set of challenges. One major concern is privacy, particularly when using third-party LLM services that require codes to be sent externally for analysis. Zhou et al. [22] emphasize the importance of using locally hosted LLMs to address these privacy concerns, although this approach has not yet been fully benchmarked for its effectiveness in static vulnerability scanning. Another critical challenge is the static nature of LLM training data. While fine-tuning techniques highlighted in [23] can improve model accuracy, LLMs still struggle to adapt to the latest security threats without frequent updates. These limitations underscore the need for integrating real-time knowledge retrieval systems to dynamically supply LLMs with up-to-date information, enhancing their detection capabilities.
III-D Enhancing LLMs with Knowledge Retrieval and SAST Integration
Recent research has explored the use of knowledge retrieval systems to address the limitations of LLMs, particularly their reliance on static training data. LLM4Vuln[24] highlights the critical role of robust knowledge retrieval in enhancing LLM-based vulnerability scanning, demonstrating that such systems can provide LLMs with current data and improve their scanning capabilities. However, this study primarily focuses on scanning smart contracts and does not deeply explore the integration of SAST results with knowledge retrieval. Du et al. [25] implement a RAG system to enhance LLMs with real-time data, using code-based similarity searches. While this approach shares similarities with our method, it does not combine functionality-similarity search or directly support SAST scanners, opting instead for a standalone approach without leveraging SAST outputs. Efforts to directly assist SAST with LLMs have been limited, with research often focusing on specific case studies rather than broader implementations. Li et al. [18] use CodeQL for detailed code querying and analysis but does not integrate SAST scanner outputs into the LLM prompt or employ advanced knowledge retrieval systems as we propose.
IV Methodology
In this section, we will first introduce the idea of combining conservative SAST scanner with LLMs in Section IV-A, and then gather vulnerability reports in Section IV-B and prepare them ready for the knowledge retrieval process in Section IV-C. Then, we design and test seven different knowledge retrieval methods in Section IV-D.
IV-A Combining conservative SAST scanners with LLMs
To implement LSAST, it is crucial to choose an appropriate SAST scanning tool. We selected the open-source SAST scanner Bearer due to its fast response times and its ability to output data in various formats, which is useful for integrating the results into the prompt for the LLM.
An example of a simplified output from a single Bearer scan finding is shown in Listing 1, with crucial details, such as the type of vulnerability and the affected lines. To better integrate these findings into the prompt for the LLM, we format the output in the way shown in Listing 2. Additionally, we need to provide the LLM with the target code we want to analyze. After importing the code file as a string, we adjust each line of the code to include line number information.
"cwe_ids": [
"78"
[…]
"title": "Unsanitized user input in OS command",
"description": "Executing operating system commands with unsanitized ...",
"line_number": 39,
"full_filename": "/Users/User/repos/dvna/core/appHandler.js",
"source": {
"start": 39,
"end": 44,
Listing 1: single Bearer scan finding (simplified)
’CWE-943 (line 59-65)’,
’CWE-943 (line 85-89)’,
’CWE-943 (line 107-111)’,
’CWE-943 (line 145-149)’,
’CWE-611 (line 235)’,
’CWE-78 (line 39-44)’
Listing 2: formated output of Bearer scan results
line 1: var db = require(’../models’)
line 2:
line 3: module.exports.userSearch = function (req, res) {
line 4: var query = "SELECT name,id FROM Users WHERE login=’" + req.body.login + "’";
line 5: db.sequelize.query(query, {
line 6: model: db.User
line 7: }).then(user => {
line 8: if (user.length) {
line 9: var output = {
Listing 3: target code with line information added
Target code string after adjustments is demonstrated in Listing 3. Including line number information in the code ensures that the SAST response, which also includes line numbers, can be accurately interpreted by the LLM. Omitting line numbers may hinder the LLM’s ability to fully utilize SAST scanner results, potentially leading to inaccuracies in the LLMs own findings regarding line-specific vulnerabilities.
Furthermore, Listing 4 presents the complete prompt, encompassing both the target code and the results from the Bearer scan. This integrated prompt, which includes detailed instructions, is now ready for direct querying by the LLM.
Code:
{target_code}
After scanning the code with a SAST Scanner we receive following vulnerabilities:
{bearer_result}
You are a very efficient vulnerability scanner.
Follow these tasks carefully:
- Evaluate whether the given code includes any more vulnerabilities other then the ones found by the SAST Scanner.
- Only output vulnerabilities that were not found by the SAST Scanner.
- There can be multiple vulnerabilities
- Only output vulnerabilites that are 100 percent certain to be present in the given code.
- If you dont find any additional vulnerabilties just answer: "no additional vulnerabilties found"
- If you find a vulnerability structure your answer like this:
"CWE-ID: <CWE-ID>
Reason: <reason why code is vulnerable>
line: <the lines that cause the vulnerability>
code-snippet: <the code-snippet that cause the vulnerability>"
Listing 4: LLM prompt with bearer results
IV-B Gathering Vulnerability Reports
To ensure our queries include the most current vulnerability information, we’ve implemented a Knowledge Retrieval System. This system aggregates vulnerability reports from the past year, offering insights into newly exploitable vulnerabilities that may not be covered in the LLMs training data.
"id": 2493548,
"title": "Incorrect Type Conversion in interpreting IPv4-mapped IPv6 addresses and below ‘curl‘ results in indeterminate SSRF vulnerabilities.",
"severity_rating": "critical",
"cve_ids": [
"CVE-2023-24329",
"CVE-2024-22243"
"cwe-id": "CWE-843",
"cwe-title": "Type Confusion",
[…]
Listing 5: HackerOne vulnerability report (simplified)
To begin, we retrieve these reports and prepare them for a similarity search. We utilize the Hacktivity API [^1] HackerOne, (https://hackerone.com/hacktivity/), a leading Bug Bounty Platform, hosts programs where companies incentivize security researchers to uncover vulnerabilities in their systems. Some findings from these programs are openly disclosed on HackerOne and accessible via their API. A collection of vulnerability reports is invaluable for LLM-supported DAST or SAST scanners, particularly in scenarios involving both Black-Box (application-only access) and White-Box (full code access) approaches. for gathering vulnerability data. Our efforts have yielded 873 vulnerability reports in JSON format, all dated within the past year and containing software code details.
We provide a simplified example of a single vulnerability report in Listing 5. This report includes the title, severity rating (ranging from low to critical), CVE IDs, CWE IDs and title, and vulnerability information. The vulnerability information section provides a detailed description of the vulnerability along with relevant code examples.
IV-C Preparing Vulnerability Reports for Knowledge Retrieval
To effectively deliver vulnerability reports to the LLM, we utilize a retrieval-augmented system that combines SAST scan results with the target code. This integration enhances the LLMs understanding by providing a more complete context for analysis. Instead of overwhelming the LLM with all 873 available vulnerability reports, which could lead to distractions and reduced performance, our approach selectively retrieves only the most relevant reports that match the specific characteristics of the code being scanned. This targeted delivery ensures that the LLM focuses on pertinent information, thereby improving its accuracy in identifying new vulnerabilities.
To determine the most relevant reports for the target code, we introduce LLM4Vuln [24], which performs a similarity search based on code functionality. The process begins by summarizing the functionality of each code snippet in the vulnerability reports and storing these summaries in a vector database using the LLM (see Listing 6). When scanning the target code (TC) for vulnerabilities, we also summarize its functionality using the LLM. We then conduct a similarity search in the vector database to find vulnerable code snippets (VC) with similar functionalities. Multiple functionalities of vulnerable code are correlated back to their respective reports using sequence numbers stored in the database, allowing us to extract and deliver the most relevant reports to the LLM.
Recognizing the weak correlation between code functionality and vulnerabilities, we developed a new approach for conducting similarity searches. A stronger correlation often emerges when examining code abstraction in relation to vulnerabilities. Abstraction, or code minimization, simplifies source code to essential components necessary for execution. To perform a similarity search based on abstraction, we extract code from vulnerability reports and apply code minification techniques using minifiers or LLM capabilities (see Listing 7). After minifying the vulnerable code snippets, we store them in a vector database.
When scanning the target code (TC) for vulnerabilities, we first minimize the TC and then search for vulnerable code (VC) with similar minified representations. We aggregate the code abstractions and correlate them back to the original reports using sequence numbers, as previously explained. This approach ensures we gather reports with structures highly resembling the target code, increasing the likelihood of identifying identical vulnerabilities present in both VC and TC.
This method closely aligns with conservative SAST scanner methodologies, which aim to identify vulnerable patterns within code by leveraging predefined rules and heuristics to detect known weaknesses, ensuring thorough and systematic analysis of potential security issues.
{target_code}
Given the code snippet, please remove all contextual information and variable names, leaving only the syntactic structure intact.
1. The task involves extracting the core syntactic structure of the code while removing all contextual information and variable names, leaving only the essential code pattern intact.
2. This process requires meticulous attention to detail and a thorough understanding of the codes structure.
3. Focus on identifying and preserving the fundamental syntax elements, such as loops, conditionals, function definitions, and method invocations, while disregarding any specific variable names, comments, or non-essential whitespace.
4. The resulting isolated code pattern will serve as a basis for further analysis and comparison, aiding in the identification of similar code patterns across multiple codebases or vulnerability reports.
5. Use the format: Code-Pattern:[xxxx], placing the extracted Code Pattern inside the brackets.
6. If there is no code only answer "Code-Pattern:[no code]"
Listing 6: LLM prompt to get abstraction of target code
{target_code}
Follow these tasks:
1. Outline the operational purpose of the provided code using imperative language, such as Calculate bank account balance.’ Your explanation should be brief, no more than one paragraph and within 40-50 words.
2. Ensure your description focuses solely on the functionality or business logic without referencing specific variables, functions, or expressions.
Listing 7: LLM prompt to get summarization of target code
IV-D Knowledge Retrieval Methods
To address the issue of LLMs relying on static training data, we use a dynamic Knowledge Retrieval System.
In this section, we provide some illustrative figures to show the various methods available for constructing the knowledge retrieval process through the combination of the functionality approach, abstraction approach, and conservative SAST scanner results.
All the notations are summarized in the Table I.
TABLE I: Definitions of Terms
Abbreviation Description
TC Target Code
SR SAST-Scanner Result
VR Vulnerability Reports
FVR Functionality-similar Vulnerability Reports
CVR Code-Abstraction-similar Vulnerability Reports
k Value of top $k$ in the similarity search
The following approaches are designed to progressively improve the LLM’s ability to identify vulnerabilities by integrating various sources of information:
IV-D1 Raw Own-Knowledge Approach (Baseline)
Figure 1: RAW Own-Knowledge Approach.
[…]
This simplest method involves directly querying the LLM using only the TC without any support from external sources or retrieval systems, as shown in Fig. 1. It relies solely on static training data, often missing newer or context-specific vulnerabilities. This approach is considered as the baseline.
IV-D2 RAW Own-Knowledge LSAST-Approach
Figure 2: RAW Own-Knowledge LSAST-Approach.
[…]
Building upon the baseline, we incorporate the SAST scanner result, SR, along with the TC into the LLM input as our raw LSAST approach, as shown in Fig. 2. This allows the LLM gain initial insights into known vulnerabilities identified by traditional static analysis to enhance the overall detection accuracy.
IV-D3 Functionality-Based Retrieval Approach
Figure 3: RAW Functionality Approach.
[…]
In this approach, depicted in Fig. 3, the LLM first summarizes the functionality of the TC. A similarity search is then conducted within a vector database containing functionalities of known vulnerable code snippets. The most relevant FVR are retrieved and provided to the LLM along with the target code.
IV-D4 Functionality LSAST-Approach
Figure 4: Functionality LSAST-Approach.
[…]
For the functionality, our approach utilizes Eq. (2) to extend the retrieval process by also incorporating SAST results (SR) alongside functionality reports (FVR) as shown in Fig. 4. Combining the detailed vulnerability reports with SAST’s known patterns provides a dual-layered context, improving the LLM’s performance by cross-referencing functionality-driven insights with existing static analysis results.
IV-D5 Abstraction-Based Retrieval Approach
Figure 5: RAW Abstraction Approach
[…]
In the abstraction approach as shown in Fig.5, the LLM abstracts the TC into its fundamental structual components. A similarity search is conducted within a vector database of code abstractions, retrieving the closest matching CVR.
IV-D6 Abstraction LSAST-approach
Figure 6: Abstraction LSAST-Approach.
[…]
As shown in Fig. 6, our abstraction LSAST-approach combines the strengths of abstraction-based retrieval with SR, providing the LLM with a comprehensive input that includes both structural patterns and known vulnerabilities. This dual approach allows the LLM to validate structural vulnerabilities against established patterns found by SAST, enhancing detection accuracy, especially in complex or poorly documented code segments.
IV-D7 Combined Approach
Figure 7: RAW Combined Approach.
[…]
This approach combines both FVR and CVR, integrating them with the target code to provide a holistic view, demonstrated in Fig. 7. k VR’s from the Functionality similarity search and k VR’s from the Pattern similarity search are gathered. These vulnerability reports are then provided to the LLM.
Figure 8: Combined LSAST-Approach.
[…]
The most comprehensive method, Combined-LSAST, integrates TC with functionality (FVR), abstraction (CVR), and bearer SAST scanner results (SR), illustrated in Fig. 8. This approach unites all previously described strategies, leveraging the strengths of each to provide a robust input that covers all angles of vulnerability detection. By combining these methods, the LLM is equipped with detailed insights that span functionality, structure, and known vulnerabilities, resulting in a significantly enhanced detection capability.
V Evaluation
Since previous work has already shown that LLMs can be effectively utilized for vulnerability scanning and it is very time-consuming to evaluate all approaches, we have decided to structure our approach in the most efficient way. First, for all evaluations, we are using Llama-3-70B, as it is a high-performing mid-sized model that can be hosted without needing very large cloud infrastructures or sending the code to third-party LLM providers. In addition, we have first evaluated which of the RAG approaches is most efficient in a small-scale evaluation before attempting the big evaluation between the raw models and the best-performing RAG approach.
V-A Comparison of Abstraction, Functionality, and Combined
We start by evaluating the Abstraction Approach (Eq. (5)), Functionality Approach (Eq. (3)), and the Combined Approach (Eq. (7)) without feeding the bearer input into the prompts given to LLM. We evaluate our approaches using four well-known open-source security testing projects that are described in the following.
  1. 1.
    DVWA (Damn Vulnerable Web Application) [^2] https://github.com/digininja/DVWA: A PHP/MySQL web application intentionally is designed with vulnerabilities to help security professionals test their tools and hone their vulnerability detection skills.
  2. 2.
    DVNA (Damn Vulnerable NodeJS Application) [^3] https://github.com/appsecco/dvna: A Node.js web application that intentionally incorporates OWASP Top 10 Vulnerabilities to demonstrate how these vulnerabilities can be identified and mitigated.
  3. 3.
    OWASP Juice Shop [^4] https://owasp.org/www-project-juice-shop/: A modern JavaScript-based application (using Node.js, Express, and Angular) intentionally built with vulnerabilities. It serves as a platform for security training and testing of security tools.
  4. 4.
    WebGoat  [^5] https://owasp.org/www-project-webgoat/: A Java-based web application deliberately including vulnerabilities, designed for security professionals to explore and understand common Java-specific security issues.
The projects include different programming languages like Java, JavaScript, TypeScript, and PHP, and are built to include a wide range of different vulnerabilities. Across all 4 projects, using our 3 scanner approaches, we have scanned a total of 10 files that include more than 15 different kinds of vulnerabilities.
TABLE II: Comparison of Abstraction, Functionality, and Combined Approaches
Measure Abstraction Functionality Combined
TP-Rate 23.81% 14.29% 28.57%
FP-Rate 100% 100% 100%
Accuracy 19.23% 10% 24%
Precision 50% 25% 60%
F1-Score 32.26% 18.18% 38.71%
For our evaluation metrics, we use standard measures in vulnerability detection:
  • True Positive Rate (TP-Rate): The proportion of reported vulnerabilities that were actual vulnerabilities.
  • False Positive Rate (FP-Rate): The proportion of vulnerabilities that were overlooked.
  • Accuracy: The ratio of correctly predicted instances to the total number of predictions.
  • Precision: The proportion of predicted vulnerabilities that were actually vulnerabilities.
  • F1-Score: A single score that represents overall performance.
We use the Bearer results as the reference standard, challenging the three approaches to identify as many vulnerabilities as the static scanner detects. This evaluation aims to determine which of the three approaches best covers the underlying vulnerabilities.
Table II illustrates that the Combined Approach has outperformed both the Abstraction Approach and the Functionality Approach. The 100% FP-Rate is achieved since all files include at least 1 vulnerability, and therefore a scan never had the opportunity to label a file as false positive.
As explained earlier, there is a higher correlation between code abstraction and underlying vulnerabilities compared to functionality and underlying vulnerabilities. The reason is that functionalities can differ greatly but still have the same vulnerabilities, and vice versa. Conservative scanners also scan for underlying code patterns or syntactic features. Therefore, we can see that our new Abstraction and Combined approaches work better than the Functionality approach, which was introduced in LLM4Vuln [24].
The effectiveness of the Combined Approach implies that the additional vulnerability reports it gathers, relative to the pattern approach, contribute valuable information to the LLM, even if they lack high similarity in code abstraction.
The overall weak performance can be explained by examining the reports that were passed to the LLM depending on the chosen approach. We have looked at the vulnerability reports that were passed to the LLM after the selected similarity search approach and found that most of the reports were not related to the CWE’s that were found by Bearer.
TABLE III: Comparison of similarity search for Abstraction and Functionality Approaches
Measure Abstraction Functionality
TP-Rate 33.33% 16.67%
FP-Rate 100% 100%
Accuracy 12.50% 6.90%
Precision 16.67% 10.53%
F1-Score 22.22% 12.90%
Table III represents the results of the Similarity Search Performance. The TP-Rate can be interpreted as how many of the vulnerabilities that we passed to the LLM were also found by Bearer. In other words, around 67% of the reports that we passed to the LLM, with the abstraction approach, were related to other vulnerabilities than the ones found by Bearer. This indicates that we effectively distracted the LLM from identifying the vulnerabilities detected by Bearer by providing a large number of unrelated vulnerabilities, highlighting the importance of refining the input to ensure the LLM remains focused on the most relevant security threats.
V-B RAW vs LSAST vs LSAST-Combined
We have evaluated that the combined approach performs the best out of the three knowledge retrieval approaches in the above section. To further explore its effectiveness, we compare the combined approach with those that do not incorporate a knowledge retrieval system. We have scanned the open-source projects DVWA, DVNA, Juice Shop, and WebGoat, using three approaches: the combined LSAST-Approach (Eq.(8)) (with knowledge retrieval and bearer results), the raw LSAST-Approach (Eq.(2)) (without knowledge retrieval, but with bearer results), and the raw Approach (Eq.(1)) (without knowledge retrieval and without bearer results). In the combined and RAW LSAST approaches, the Bearer results are directly provided to the LLM. We aimed to determine how many true vulnerabilities the LLM could identify beyond those already detected by Bearer. To classify the vulnerabilities found by the scanners, we manually review each detected vulnerability and labeled them as follows:
  • True Positive: If the predicted vulnerability is labeled by us as a real vulnerability that has not been found by Bearer.
  • False Positive: If the predicted vulnerability is not a real vulnerability that can be exploited or is a hallucination.
  • False Negative: If one of the other approaches has found a real vulnerability but they have not been found by this approach.
Vulnerabilities already detected by Bearer are marked as duplicates and excluded from the evaluation.
TABLE IV: Comparison of Measures.
Measure raw raw-LSAST Combined-LSAST
TP-Rate 17.91% 68.89% 35.71%
FP-Rate 100% 100% 100%
Accuracy 24% 62% 32.61%
Precision 63.16% 86.11% 78.95%
F1-Score 38.71% 76.54% 49.18%
The results in Table IV indicate that including Bearer findings significantly enhances the LLM’s ability to detect vulnerabilities that static scanners miss. However, the RAW LSAST approach outperforms the Combined LSAST approach, while the RAW approach yields the lowest performance.
Two key insights emerge from our evaluation: first, providing SAST scanner results to the LLM enhances its capability to identify new vulnerabilities that cannot be detected by SAST scanners alone. Second, a knowledge retrieval framework that delivers similar vulnerabilities to the LLM can negatively impact performance if not carefully configured.
Integrating SAST scanner results help the LLM focus on detecting vulnerabilities in code areas not previously identified by Bearer, thereby improving overall detection accuracy. Conversely, our findings indicate that using a RAG infrastructure can be counterproductive. If the vulnerability reports provided to the LLM lack high relevance to the analyzed code segments, they may divert the LLM’s attention and reduce its effectiveness.
V-C Real-world test
To demonstrate that LSAST is capable of identifying vulnerabilities missed by conventional SAST scanners in real-world scenarios, we scan the Python package pyDust  [^6] PyDust (https://github.com/MondoAurora/pydust/tree/main/dust) is a complex Python package that implements an entity-attribute-value data model, providing mechanisms for creating, manipulating, and serializing entities with dynamic attributes, along with features for change tracking and artifact management. for vulnerabilities and examine the findings in collaboration with the repository’s authors.
After scanning the repository, LSAST correctly identifies 2 High-Risk, 1 Medium-Risk, and 3 Low-Risk vulnerabilities that conventional scanners failed to detect. These vulnerabilities could lead to serious security issues, such as code injection or remote code execution in the application if left unfixed.
In addition to the 5 true positives, the scanner also reports 2 false positives, highlighting the importance of double-checking scanner results.
Given that PyDust is a sophisticated piece of software and LSAST is still able to find vulnerabilities, we have demonstrated that LSAST can perform effectively in real-world environments.
VI Conclusion
This study introduces LSAST, integrating LLMs with static application security testing to enhance vulnerability detection. By combining SAST results and a knowledge retrieval system, LSAST improves the detection of complex vulnerabilities beyond the capabilities of traditional scanners.
Our results show that integrating SAST with LLMs significantly boosts detection accuracy, though the performance of knowledge retrieval depends heavily on the relevance of the data provided. The research addresses key challenges, including privacy concerns and the limitations of static training data, contributing to more effective cybersecurity analysis. Moreover, the real-world test on the Python package pyDust validates the practical applicability of LSAST, showing its capability to identify high-risk vulnerabilities that conventional scanners miss.
Future work will refine knowledge retrieval strategies and explore deeper integration of LLMs with dynamic testing methods to further strengthen vulnerability detection systems. Another promising direction for future research is the development of adaptive learning mechanisms that allow LLMs to continuously update their knowledge base with the latest security threats and mitigation techniques.